OpenStudy (anonymous):

Solve the system by graphing. 2x + y = 6 4x + 2y = 12

OpenStudy (campbell_st):

well before you graph the lines... take the 2nd line and divide every term by 2, what do you get...?

OpenStudy (anonymous):

2x+y=12

OpenStudy (anonymous):

2x+y=6

OpenStudy (campbell_st):

nearly... you need to halve every term... including the 12... so what do you get..?

OpenStudy (campbell_st):

great... so what can you say about the 2 lines... before graphing...?

OpenStudy (anonymous):

That they are the same.

OpenStudy (campbell_st):

yeap... so you only need to graph 1 line... and the have an infinite numbers of solutions or points of intersection

OpenStudy (anonymous):

Thanks :D C) {(x, y)| 2x + y = 6}

OpenStudy (jmark):

Given 2x + y = 6.............(1) 4x + 2y = 12 …..............(2) divide equation (2) by 2 =>2x+y=6 this is same as the equation (1) hence there is only one equation =>y=-2x+6 where slope m=-2 and c=6 See also CBSE Sample Paper http://goo.gl/TqhFe1